OpenStudy (anonymous):

Use the disk or shell method to find the volume of the solid generated by revolving the region in the first quadrant bounded by the graph of the equation about the given line. (i) the x-axis (ii) the y-axis

OpenStudy (anonymous):

\[x^\frac{ 2 }{ 3 }+y^\frac{ 2 }{ 3 }=11^\frac{ 2 }{ 3 }\]

zepdrix (zepdrix):

Ooo that's an interesting shape! :o

zepdrix (zepdrix):

https://www.desmos.com/calculator/wbefoatq8z

zepdrix (zepdrix):

So should we try spinning around the x-axis first maybe?